OpenStudy (anonymous):

Can someone help me? Choose the correct slope of the line that passes through the points (1, −3) and (3, −5). -1 1 0 1/2

OpenStudy (astrophysics):

You may apply the slope formula \[ m = \frac{ y_{2}-y _{1} }{ x _{2}-x _{1} } \]

OpenStudy (astrophysics):

\[(x_1,y_1)~~~\text{and}~~~(x_2,y_2)\] are your points

OpenStudy (anonymous):

@Astrophysics \[-5--3/3-1 \]

pooja195 (pooja195):

Good start.

OpenStudy (anonymous):

That's all I know how to do @pooja195 I haven't done this since 8th grade lol.

pooja195 (pooja195):

Ok when you have 2 negative signs together they become a positive sign.So we would ADD. -5+3=?

OpenStudy (anonymous):

-2 @pooja195

pooja195 (pooja195):

Good :) 3-1=?

OpenStudy (anonymous):

2 @pooja195

pooja195 (pooja195):

Good :) Slopes can be simplified so in this case divide: -2/2=?

pooja195 (pooja195):

*Fractions

OpenStudy (anonymous):

0? @pooja195

OpenStudy (anonymous):

:)

pooja195 (pooja195):

Nope we are dividing

OpenStudy (anonymous):

\[\frac{ -2 }{ 2 }\]

OpenStudy (anonymous):

ohh -1 @pooja195

pooja195 (pooja195):

Correct! :D

OpenStudy (anonymous):

Nice!!! Very Nice!

OpenStudy (anonymous):

May I ask you a few more @pooja195

pooja195 (pooja195):

Go for it

OpenStudy (anonymous):

Given the equation y − 4 = three fourths(x + 8) in point-slope form, identify the equation of the same line in standard form. -3/4x+y=10 3x-4y=-40 y=3/4x+12 y=3/4xx+10

OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (−1, 6) and has a slope of −3. y-6=-3x-3 y-6=-3(x+1) y=-3x+3 3x+y=3 @pooja195

pooja195 (pooja195):

can you post one at a time i dont do well with standered form hopefully @Astrophysics can help with that one

OpenStudy (anonymous):

Sorry. I haven't done any of this in about 4 years lol

OpenStudy (astrophysics):

\[y-4=\frac{ 3 }{ 4 }(x+8)\]

OpenStudy (astrophysics):

They want it in standard form eh, that's \[Ax+By=C\] in this case anyways, a little algebra

OpenStudy (anonymous):

Um how would I go about the fraction? @Astrophysics

OpenStudy (astrophysics):

 can you finish it off